OpenStudy (anonymous):

3x+2y=19 3x+3y=-6 solve the simultaneous equation by using elimination method.

OpenStudy (anonymous):

3x+2y=19 3x+3y=-6--> Multiply by -1: 3x+2y=19 -1(3x+3y=-6) 3x+2y=19 -3x-3y=6 ----------- -1y=25 y= -25 Now, plug it back into the first equation: 3x+2(-25)=19 3x-50=19 3x=69 x=23 x=23, y= -25 Does it make sense? :o)

OpenStudy (anonymous):

@girleegurl In this case should you add or subtract using elimination method?